OpenStudy (anonymous):

For a positive integer n, let Pn denote the product of the digits of n, and Sn denote the sum of the digits of n. The number of integers between 10 and 1000 for which Pn + Sn = n is

OpenStudy (perl):

did you try some numbers

OpenStudy (perl):

P(10)=1*0 = 0 S(10) = 1 + 0 = 1 P(11) = 1*1 = 1 S(11) = 1 + 1 = 2

OpenStudy (ikram002p):

so \(\large n=a_0 10^0 +a_110^1 +a_210^2 +...+a_m 10^m = \sum_{i=0}^m a_i\times 10^i\\ \large P_n=a_0\times a_1 \times a_2...\times a_m=\prod _{i=0}^m a_i \\ \large S_n=a_0+ a_1 + a_2...+ a_m=\sum _{i=0}^m a_i\)

OpenStudy (ikram002p):

hehe i should try code it :)

OpenStudy (perl):

yes to program this would work . i dont see an obvious algebraic trick

OpenStudy (ikram002p):

idk xD i'll see also programming might give a pattern right ?

OpenStudy (perl):

it might , or just count up the number of answers

ganeshie8 (ganeshie8):

\[ \prod _{i=0}^m a_i + \sum _{i=0}^m a_i=\sum_{i=0}^m a_i 10^i \]

OpenStudy (ikram002p):

ok then we should expand ?

ganeshie8 (ganeshie8):

idk i have just copy pasted that from ur reply abvoe

OpenStudy (ikram002p):

ik xD i thought u might have sme idea

ganeshie8 (ganeshie8):

this looks more like a program problem as perl said

ganeshie8 (ganeshie8):

@dan815

ganeshie8 (ganeshie8):

notice that m is fixed here : m=2

ganeshie8 (ganeshie8):

they just want the solutions for 3 digit numbers

ganeshie8 (ganeshie8):

abc + (a+b+c) = 100a + 10b + c

ganeshie8 (ganeshie8):

thats a number theory problem

OpenStudy (perl):

should we include 1000 as well, we can check that one manually

OpenStudy (perl):

its false for 1000

OpenStudy (ikram002p):

ok here is my algorithm 1- set numbers from 10 to 100 in array as n[990] 2- send n[i] to Pn and Sn function 3- if statement to check

OpenStudy (anonymous):

So much smart people. XD

ganeshie8 (ganeshie8):

abc + (a+b+c) = 100a + 10b + c 99a + 9b = abc 11a + b = abc/9 this is just a diophantine equation in "two" variables solve the equation for `11a+b = 1` and extrapolate the results

OpenStudy (ikram002p):

oh lol abc+a+b+c =100 a +10 b +c

OpenStudy (ikram002p):

ok lets do it :)

ganeshie8 (ganeshie8):

11a + b = 1 by inspection : a = 1 b = -10 is one solution, so all other solutions can be given by : `a = 1+t`, `b=-10-11t`

ganeshie8 (ganeshie8):

next, we can find the values of c, such that `abc/9` is an integer

ganeshie8 (ganeshie8):

still wid me ikram

OpenStudy (perl):

why did you set it equal to 1 11a + b = 1

OpenStudy (ikram002p):

yepp

ganeshie8 (ganeshie8):

that allows us to multiply any number both sides and solve for exampile, we can multiply abc/9 both sides here

ganeshie8 (ganeshie8):

if x,y are solutions of 11a + b = 1 then x*abc/9 , y*abc/9 will be solutions of 11m + n = abc/9 yeah this is just not good enough

OpenStudy (ikram002p):

but still give an idea like conditions :O

ganeshie8 (ganeshie8):

because we want to solve 11a+b = abc/9 and not 11m+n = abc/9

OpenStudy (ikram002p):

i got ur point hmm

ganeshie8 (ganeshie8):

anyways atleas this much is clear : 9 | abc

ganeshie8 (ganeshie8):

remember the divisibility rule for 9 ?

OpenStudy (ikram002p):

so one of them must be 9 , since they are unit digits

ganeshie8 (ganeshie8):

oh right right !

ganeshie8 (ganeshie8):

one of them has to be 9

OpenStudy (perl):

if the sum of digits is divisible by 9

OpenStudy (ikram002p):

wait , a=9 or b=9 or c=9 or a=3 and b=3 something like this

ganeshie8 (ganeshie8):

its a*b*c not abc, my bad im fully on offtrack on this problem

ganeshie8 (ganeshie8):

yup!

ganeshie8 (ganeshie8):

9 | a*b*c

ganeshie8 (ganeshie8):

thats just an observation, wont help us solve the problem

ganeshie8 (ganeshie8):

here is the answer if it helps you revere enginner ``` 19 29 39 49 59 69 79 89 99 ```

OpenStudy (ikram002p):

how did u get it ?

ganeshie8 (ganeshie8):

program

OpenStudy (ikram002p):

xD mine dint work ;_;

ganeshie8 (ganeshie8):

here is mine ``` foreach my $n (11 .. 999) { my @digits = split //, $n; my $sum = eval join '+', @digits; my $prod = eval join '*', @digits; print "$n\n" if ($sum+$prod == $n); } ```

ganeshie8 (ganeshie8):

but this problem should be solved using number theory, the solutios are only few and they look pretty patterny

OpenStudy (ikram002p):

nvm xD im not good in programming

OpenStudy (ikram002p):

yeah :)

OpenStudy (ikram002p):

so basically , 9|abc

ganeshie8 (ganeshie8):

9|a*b*c

OpenStudy (ikram002p):

yeah i now , if im gonna refer to n i would use 100a+10b+c mathematician xD

ganeshie8 (ganeshie8):

okie

ganeshie8 (ganeshie8):

from the output, a = 0 any idea why it has to be so ?

OpenStudy (ikram002p):

so let ok i might have an idea

OpenStudy (ikram002p):

9*9*9=729 9+9+9=27 so 729+27=756 so a cant be 9 :)

ganeshie8 (ganeshie8):

a=0 gives u 2 digit numbers, so we should consider 2 digit numbers and 3 digit numbers separately hmm

OpenStudy (ikram002p):

hmm 100a+10b+c = abc + a+b+c so any digit we would take for a 100a+10b+c > abc + a+b+c

OpenStudy (ikram002p):

:D thus we consider a=0

OpenStudy (ikram002p):

10b+c=bc+b+c

ganeshie8 (ganeshie8):

ohk keep going

OpenStudy (ikram002p):

9b=bc Xd c=9

ganeshie8 (ganeshie8):

wow! that was easy

OpenStudy (ikram002p):

so all 19 29 39 49 59 69 79 89 99

ganeshie8 (ganeshie8):

nice :)

OpenStudy (ikram002p):

yeah i guess the trick in making a=0

ganeshie8 (ganeshie8):

im thinking about this : 100a+10b+c = abc + a+b+c so any digit we would take for a 100a+10b+c > abc + a+b+c is there a proof for this

OpenStudy (ikram002p):

lol i wont note that without seeing output xD

OpenStudy (ikram002p):

yeah let proof it

OpenStudy (ikram002p):

99a+9b>ab

ganeshie8 (ganeshie8):

claim : the value of a positive integer is always greater than the product of its digits

OpenStudy (ikram002p):

highest value that ab can take is 9*9 < 100

ganeshie8 (ganeshie8):

however the value of n is 10b + c only

OpenStudy (ikram002p):

:D

ganeshie8 (ganeshie8):

no its not clear to me yet

ganeshie8 (ganeshie8):

we want to prove 100a + 10b + c > a*b*c

OpenStudy (ikram002p):

oh lol ok 100a+10b+c <abc+a+c+b 99a+9b<abc now highest value that abc could take is 9*9*9 right ?

ganeshie8 (ganeshie8):

yes

OpenStudy (ikram002p):

so in other way we wanna prove this 99a+b > 9*9*9 99*9+9 > 9*9*9 true

OpenStudy (ikram002p):

wait hmm

ganeshie8 (ganeshie8):

100a+10b+c = 10(10a + b) + c > b(ca + b) = abc + b^2 > abc

ganeshie8 (ganeshie8):

xD

ganeshie8 (ganeshie8):

got it :) that was a very useful trick to stop after 2 digit numbers

OpenStudy (ikram002p):

wait what u did :O

ganeshie8 (ganeshie8):

100a+10b+c = 10(10a + b) + c

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (ikram002p):

yeah keep going

ganeshie8 (ganeshie8):

10 > b and 10 > c so : 100a+10b+c = 10(10a + b) + c > b(ca + b) + c

ganeshie8 (ganeshie8):

fine ?

OpenStudy (ikram002p):

neat xD yeah fine

ganeshie8 (ganeshie8):

i have replaced 10's with b and c and put the inequality

ganeshie8 (ganeshie8):

100a+10b+c = 10(10a + b) + c > b(ca + b) + c = abc + b^2 + c > abc

OpenStudy (ikram002p):

nishe xD

OpenStudy (ikram002p):

hehe i tried to show it with maximum bhaha by taking 9 as highest blah but this is more neat